OpenStudy (anonymous):

\[\int\frac{ (lnx) ^{2} }{ x }dx\]

OpenStudy (anonymous):

u= 2ln(x) du = 2/x

OpenStudy (anonymous):

then

OpenStudy (anonymous):

integral of 1/2*u*du

OpenStudy (anonymous):

and

OpenStudy (anonymous):

u srs?

OpenStudy (anonymous):

or should I say u^2/2 srs?

OpenStudy (anonymous):

don't get it, because the answer is \[\frac{ 1 }{3 }lnx ^{3}+c\]

OpenStudy (anonymous):

i just don't know how to solve it

OpenStudy (anonymous):

don't think you can write it like that...

OpenStudy (calculusfunctions):

Alright, do you know the substitution rule for integrals? Also do you know the derivatives of logarithm functions? If you do, then we're off to a great start!

OpenStudy (anonymous):

yep

OpenStudy (calculusfunctions):

Were you replying to me or Algebraic?

OpenStudy (anonymous):

you

OpenStudy (calculusfunctions):

OK, so then when you see\[\int\limits_{}^{}\frac{ \ln x ^{2} }{ x }dx\]What do you see as the first potential step?

OpenStudy (callisto):

Sorry to interrupt, is the question (i) \(\int \frac{lnx^2}{x}dx\) or (ii) \(\int \frac{(lnx)^2}{x}dx\)? They are different...

OpenStudy (calculusfunctions):

Do you know your logarithm properties? for example\[\ln x ^{n}=n \ln x\]Do you know this property?

OpenStudy (anonymous):

yeah it's supposed to be (ln(x))^2

OpenStudy (calculusfunctions):

@Callisto, it's the former.

OpenStudy (callisto):

For the first one, I don't think you can get (1/3)(lnx)^3 +C

OpenStudy (anonymous):

nope:)

OpenStudy (anonymous):

turn it to \[\int\limits_{?}^{?}\frac{ 1 }{ x } \times \ln x ^{2} dx\]

OpenStudy (anonymous):

that clear it up @lambchamps ?

OpenStudy (anonymous):

the second @Callisto

OpenStudy (callisto):

You see... That's why...

OpenStudy (anonymous):

sorry it is the second guys

OpenStudy (anonymous):

so it's u = ln(x) u^2 = (ln(x))^2 du = 1/x go to town.

OpenStudy (calculusfunctions):

@Callisto, you are right but @lambchamps, Is the question written correctly before we proceed further?

OpenStudy (anonymous):

@calculusfunctions it's the second. based on what @Callisto have mentioned

OpenStudy (anonymous):

@Algebraic! please continue

OpenStudy (calculusfunctions):

OK! so then we don't need the logarithm property I proposed earlier.

OpenStudy (anonymous):

that's it man, plug em in and integrate.

OpenStudy (calculusfunctions):

@lambchamps, do you now what the derivative of\[y =\ln x\]is?

OpenStudy (anonymous):

@calculusfunctions he or she is in calc. 2 so yeah probably. good question though.

OpenStudy (anonymous):

dy = 1/x dx?

OpenStudy (calculusfunctions):

Great, so then if\[\int\limits_{}^{}\frac{ (\ln x)^{2} }{ x }dx\]then in order to apply the substitution rule what should u equal?

OpenStudy (anonymous):

\[=\int\limits_{}^{}\frac{ 1 }{ x } \times \ln x ^{2} dx\]

OpenStudy (anonymous):

i mean (ln x)^2

OpenStudy (calculusfunctions):

First of all you keep confusing the issue by writing\[\ln x ^{2}\]instead of\[(\ln x)^{2}\]They are not the same!\[(\ln x)^{2}\neq \ln x ^{2}\]Do you understand? So we're not going to get anywhere until this gaffe is resolved.

OpenStudy (calculusfunctions):

OH, OK! Sorry, I see you did fix it.

OpenStudy (anonymous):

sorry it's \[(\ln x)^{2}\]

OpenStudy (calculusfunctions):

Do you notice that the derivative of ln x is in the integrand? So then what should u equal?

OpenStudy (anonymous):

don't know please do tell

OpenStudy (anonymous):

is it the \[\frac{ (\ln x)^{n+1} }{ n+1 } ?\]

OpenStudy (calculusfunctions):

Here are your options. Do you think u should equal a). ln x or b). 1/x c). (ln x)²

OpenStudy (calculusfunctions):

Which option do you think? a, b, or c? Just keep in mind that whichever option you choose, it's derivative must be in the integrand.

OpenStudy (anonymous):

the derivative of ln x is 1/x so i choose b

OpenStudy (calculusfunctions):

NO! I said that the derivative of the chose option must be in the integrand. NOT the antiderivative of the option must be in the integrand.

OpenStudy (calculusfunctions):

So what should u equal?

OpenStudy (anonymous):

a?

OpenStudy (calculusfunctions):

Excellent!

OpenStudy (anonymous):

and then

OpenStudy (calculusfunctions):

So now\[u =\ln x\]so then\[du =?\]

OpenStudy (anonymous):

1/x

OpenStudy (calculusfunctions):

\[du =\frac{ 1 }{ x }dx\]OK?

OpenStudy (anonymous):

ok sorry

OpenStudy (calculusfunctions):

So now if you substitute u and du into your integral, what do you have?

OpenStudy (calculusfunctions):

@integralsabiti, how is giving the solution helping the student who is trying to learn? I spent all this time trying to teach @lambchamps so that she can then do other similar problems with confidence, and you just came in wasted her time and my effort. NOT COOL!

OpenStudy (anonymous):

@calculusfunctions effort appreciated.. thanks to both of you

OpenStudy (anonymous):

sorry for interrupting .you may go on

OpenStudy (calculusfunctions):

@integralsabiti, thank you, no worries now that I know your intentions were genuine.

OpenStudy (calculusfunctions):

@lambchamps, are you still there? I'm still waiting for your response to my last question regarding your problem.

OpenStudy (anonymous):

so I would need to find the dx then substitute the value to it?

OpenStudy (anonymous):

ok calculausfunction after your done here would you help me on my problem pleases

OpenStudy (calculusfunctions):

So what do you you now have after substituting\[u =\ln x\]and\[du =\frac{ 1 }{ x }dx\]into\[\int\limits_{}^{}\frac{ (\ln x)^{2} }{ x }dx\]

OpenStudy (calculusfunctions):

Of course but let's hurry and finish this one first because I have to log out soon.

OpenStudy (calculusfunctions):

So what does the integral look like after substitution?

OpenStudy (anonymous):

dx would be xdu?

OpenStudy (calculusfunctions):

NO! If \[u =\ln x\]and\[du =\frac{ 1 }{ x }dx\]then\[\int\limits\limits_{}^{}\frac{ (\ln x)^{2} }{ x }dx =\int\limits_{}^{}u ^{2}du\]Do you see how?

OpenStudy (anonymous):

so the 1/x in (ln x)^2/x would be cancelled?

OpenStudy (calculusfunctions):

@mikala, I'll help you right after I finish here.

OpenStudy (anonymous):

because of the x.du

OpenStudy (anonymous):

please tell me i'm right

OpenStudy (calculusfunctions):

No, the\[\frac{ 1 }{ x }dx\]is being replaced with du because\[du =\frac{ 1 }{ x }dx\]Please tell me you see that.

OpenStudy (calculusfunctions):

Cancelled is a poor choice of words.

OpenStudy (anonymous):

ok, but where did the 1/x go? the one below (ln x)^2

OpenStudy (anonymous):

oh i'm sorry didn't see at first

OpenStudy (calculusfunctions):

\[\int\limits\limits_{}^{}\frac{ (\ln x)^{2} }{ x }dx\]is exactly the same as\[\int\limits_{}^{}(\frac{ 1 }{ x }∙(\ln x)^{2})dx\]correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i see the 1/x dx

OpenStudy (anonymous):

after finding that, then is the time to integrate, right?

OpenStudy (calculusfunctions):

So I replaced the factor of\[\frac{ 1 }{ x }dx\]with du because\[du =\frac{ 1 }{ x }dx Yes, now you may integrate\[\int\limits_{}^{}u ^{2}du\]

OpenStudy (calculusfunctions):

Sorry, I don't know what happened there let me try again.

OpenStudy (anonymous):

ok

OpenStudy (calculusfunctions):

Yes, now find the\[\int\limits_{}^{}u ^{2}du\]Can you do that please?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }u ^{3} + c\]

OpenStudy (calculusfunctions):

Right!

OpenStudy (calculusfunctions):

Now what's the final step?

OpenStudy (anonymous):

then substitute

OpenStudy (anonymous):

ln x right?

OpenStudy (calculusfunctions):

Yes so can you please write the final answer now?

OpenStudy (anonymous):

\[\frac{ (\ln x)^{3} }{ 3 } + C\] ayt?

OpenStudy (calculusfunctions):

Perfect!!!

OpenStudy (anonymous):

on more question is the c suppose to capitalized?

OpenStudy (anonymous):

one*

OpenStudy (calculusfunctions):

That doesn't matter. C represents a constant. Whether you write c or C or k or K etc. is irrelevant. Just don't use x, y, or z.

OpenStudy (calculusfunctions):

The most commonly used ones are c and k.